can you please explain this question with detailed explanation and background theory both?
27 6:34 119 5928 51% chapter_12_powerpoint .ppt What is the degree of unsaturation of C7H6BrNO? A. 3 B.4 4.5 5.6 E. 7
Answer #1

Degree of unsaturation:

DOU = (2C + 2 + N - X - H)/2

X is number of halogens

C is number of carbon

N is number of Nitrogen

H is number of Hydrogen

DOU = (2C + 2 + N - X - H)/2

= (2*7 + 2 + 1 - 1 - 6)/2

= (17 - 7) / 2

= 10/2

= 5

Answer: 5